OpenStudy (anonymous):

does anybody know how to write this ellipse in standard form. They give you the foci (1,0) & (1,4) and a major axis of length 6

OpenStudy (amistre64):

it really does take a bit for this information ....

OpenStudy (amistre64):

a drawing helps

OpenStudy (amistre64):

if we go ahead and center this at the origin, we can move it back later a length of 6 would amount to it crossing at (0,+- 3) \[\frac{x^2}{}+\frac{y^2}{3^2}=1\] the other parts needs to be worked as a triangle

OpenStudy (amistre64):